# I-SHEEP: Self-Alignment of LLM from Scratch through an Iterative Self-Enhancement Paradigm

### v2

@@ Results
Table 2 presents the chat benchmark performance for the Qwen-1.5 72B model across various iteration settings. More benchmark results are available in Appendix C C. Our findings are as follows: (1) Training the base model with data from the last iteration model is effective for iterative self-enhancement. At the third iteration in the One_base Setting, training the base model with the last iteration data achieves the highest performance on the chat benchmark. The notable performance improvement under this setting suggests that the model has the potential for further enhancement (refer to Table 1 72B results). Therefore, we chose the One_base setting for all subsequent experiments. (2) The data size is not the main factor influencing iterative improvement. Training the base model with the last iteration data at the 3rd iteration outperforms training the base model with a combination of all data from previous iterations.
Threshold Ablation
@@
Data Size Ablation
@@
Metacognitive Self-Assessment Analysis
Self-Assessment Robustness Analysis
Table 3 shows the performance of various self-assessment degrees in the first three iterations. See the Appendix CC for more benchmark results. The following findings can be drawn from the table: (1) Using explicit self-assessment prompt is better than using simple model internal states. On all four IFEval accuracies, the highest values are obtained in the setting where the model is explicitly prompted for self-assessment. (2) The I-SHEEP framework is robust to prompt. Although the criteria differ between simple and combined standard prompt settings, their performance is quite similar. Even without designing a prompt, using just a few examples for ICL can achieve comparable results.
@@
Self-Assessment Level Analysis.
@@ Conclusion
Limitation and Future Work
While the I-SHEEP framework can enhance model performance, the extent of final improvement after the RLHF phase remains uncertain. The complete self-improvement process (SFT+RLHF) needs further investigation, which we leave to future work. Additionally, there are increasing ethical concerns about using synthetic data, as it may intensify biases and harmful content in model responses. Although this paper employs strict filtering for generated data to reduce incorrect cognition, it cannot eliminate them. This issue also persists in models aligned with human-annotated data. Furthermore, the prompt evaluation standards we propose can be replaced with safety-related standards to enhance model safety, similar to Dromedary (Sun et al. 2023b).
Appendix A Appendix A – Self-Assessment Prompt Content
@@
Appendix B Appendix B – Case Study
@@
Appendix C Appendix C – More Benchmark Results Evaluated by Opencompass
@@
Appendix D Appendix D – MT-Bench
| single turn score | coding | extraction | humanities | math | reasoning | role play | stem | writing | total average | ||
@@ Appendix D Appendix D – MT-Bench
Appendix E Appendix E –Lora Hyperparameters and LLaMA Factory Template
@@
Appendix F Appendix F – Data quality analysis across different iterations

### v3

@@ 5.2 Iterative Setting Results
Table 2 presents the chat benchmark performance for the Qwen-1.5 72B model across various iteration settings. More benchmark results are available in Appendix C. Our findings are as follows: (1) Training the base model with data from the last iteration model is effective for iterative self-enhancement. At the third iteration in the One_base Setting, training the base model with the last iteration data achieves the highest performance on the chat benchmark. The notable performance improvement under this setting suggests that the model has the potential for further enhancement (refer to Table 1 72B results). Therefore, we chose the One_base setting for all subsequent experiments. (2) The data size is not the main factor influencing iterative improvement. Training the base model with the last iteration data at the 3rd iteration outperforms training the base model with a combination of all data from previous iterations.
5.3 Threshold Ablation
@@
5.4 Data Size Ablation
@@
5.5 Metacognitive Self-Assessment Analysis
5.5.1 Self-Assessment Robustness Analysis
Table 3 shows the performance of various self-assessment degrees in the first three iterations. See the Appendix C for more benchmark results. The following findings can be drawn from the table: (1) Using explicit self-assessment prompt is better than using simple model internal states. On all four IFEval accuracies, the highest values are obtained in the setting where the model is explicitly prompted for self-assessment. (2) The I-SHEEP framework is robust to prompt. Although the criteria differ between simple and combined standard prompt settings, their performance is quite similar. Even without designing a prompt, using just a few examples for ICL can achieve comparable results.
@@ 5.5.1 Self-Assessment Robustness Analysis
5.5.2 Self-Assessment Level Analysis.
@@ 6 Conclusion
7 Limitations
While the I-SHEEP framework can enhance model performance, the extent of final improvement after the RLHF phase remains uncertain. The complete self-improvement process (SFT+RLHF) needs further investigation, which we leave to future work. Additionally, there are increasing ethical concerns about using synthetic data, as it may intensify biases and harmful content in model responses. Although this paper employs strict filtering for generated data to reduce incorrect cognition, it cannot eliminate them.
Appendix A Self-Assessment Prompt Content
@@
Appendix B Case Study
@@
Appendix C More Benchmark Results Evaluated by Opencompass
@@
Appendix D MT-Bench
| single turn score | coding | extraction | humanities | math | reasoning | role play | stem | writing | average | ||
@@ Appendix D MT-Bench
Appendix E Lora Hyperparameters and LLaMA Factory Template
@@
Appendix F Data quality analysis across various iterations